To resolve a capacity bottleneck in its manufacturing operations, McGuire Machinery has decided to add new equipment. Vendors for two reputable firms were requested to submit proposals. The fixed costs for proposal A are $60,000, and for proposal B, $75,000. The variable cost for A is $14.00, and for B, $11.50. The revenue generated by each unit is $25.

a) What is the break-even point in units for proposal A?

b) What is the break-even point in units for proposal B?

c) Which vendor should be chosen? Why?

Please show work.
Answer #1

Break even point = Fixed Cost / ( Revenue - Variable Cost)

A. BEP for proposal A

= 60,000 / ( 25 - 14) = 5,454.54

B. BEP for proposal B

= 75,000 / ( 25 - 11.5) = 5,555.55

C. Vendor should choose the proposal A as the break even point of proposal A is lower than proposal B.